OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 10 meters and legs that measure 6 meters and 8 meters. What is the surface area of this prism if the height of the prism is 8 meters?

OpenStudy (anonymous):

curved surf. area=288 total surf. area=236

Directrix (directrix):

SA = ph + 2B where p is the perimeter of the triangular base and B is the area of the triangular base. SA = (8+6+10)*8 + 2 ((1/2)*6*8) SA = 24*8 + 48 SA = 192 + 48 SA = 240 square meters